# SALMONN: Towards Generic Hearing Abilities for Large Language Models

### v1

@@ Appendix B Changes in Perplexity during Training
Appendix C Calculation of following rate (FR) or accuracy (Acc) for SQQA, SF, Story and SAC tasks
@@
Appendix D Prompt Template for SALMONN
To train SALMONN generate responses to a text instruction given an audio input, we utilize a prompt template as follows:
@@ Appendix D Prompt Template for SALMONN
Appendix E Prompts Interacting with GPT3.5
In this work, we utilize GPT3.5 to help us generate some data and evaluate the quality of the model output automatically. We list our prompts for different purposes in Table 6. Words with all capital letters in the prompt will be replaced with the appropriate description before fed into GPT3.5.
@@
To evaluate whether the model attempts to do the speech audio coreasoning (SAC) task.
@@
To evaluate whether the model successfully complete the SAC task.

### v2

@@ Appendix D Changes in Perplexity during Training
Appendix E Calculation of following rate (FR) or accuracy (Acc) for SQQA, SF, Story and SAC tasks
@@
Appendix F Prompt Template for SALMONN
To train SALMONN to generate responses to a text instruction given an audio input, we utilize a prompt template as follows:
@@ Appendix F Prompt Template for SALMONN
Appendix G Prompts Interacting with GPT3.5
In this work, we utilize GPT3.5 to help us generate some data and evaluate the quality of the model output automatically. We list our prompts for different purposes in Table 6. Words with all capital letters in the prompt will be replaced with the appropriate description before being fed into GPT3.5.
Appendix H Potential to Extend SALMONN to Speech and Audio Generation
Although SALMONN is designed to focus on enabling LLMs with hearing abilities, it is possible to extend SALMONN to speech generation. [^3] Added at the request of an anonymous reviewer. We thank the reviewer for the suggestion. The human speech production mechanism is related to auditory perception. A well-known phenomenon attributed to the speech chain is the “Lombard reflex” which describes the effect where individuals raise their voice level to be heard more clearly while speaking in noisy environments (Lane & Tranel, 1971). This reveals the importance of having generic hearing abilities when empowering AI with fully human-like speaking abilities, and the opportunities in text-to-speech (TTS) synthesis enabled by SALMONN. This also matches the recent development in TTS that the text and audio contexts from the surrounding utterances are useful to achieve more natural prosody modelling and enable the use of more natural and casual speech data (Xu et al., 2021; Guo et al., 2021; Oplustil-Gallegos et al., 2021; Zhang et al., 2023c).
@@ Appendix H Potential to Extend SALMONN to Speech and Audio Generation
To evaluate whether the model attempts to do the speech audio co-reasoning (SAC) task.
@@
To evaluate whether the model successfully completes the SAC task.